OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL Danielle went ice skating.It cost $3 an hour and $5 for skate rental. If Danielle spent $11, how many hours did she ice-skate? Let h represent the number of hours spent ice-skating. Identify the correct equation to represent this problem and the correct solution. A. 3h - 5 =11; h = 2 B. 5h +3 = 11; h =1.6 C. 3h +5 = 11 ; h =2 D. 5h - 3 = 11 ; h =2.8

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

If you want me to explain more, then just ask. 3*h(2)+5=11 3*2=6 6+5=11, therefore C is correct.

OpenStudy (anonymous):

OH OK